# Bridging Fairness and Explainability: Can Input-Based Explanations Promote Fairness in Hate Speech Detection?

### v1

@@ Results
The results in Figures 2 and 3 highlight the limitations of using explanations for model selection. Although some methods occasionally show high rank correlations (e.g., Grad L2 for race and religion biases in BERT and Occlusion-based methods for gender and religion biases in Qwen3), none of them consistently reach the baseline of using the individual unfairness on the validation set. This limitation is particularly evident in decoder-only models, where the baseline achieves a perfect rank correlation of 1. Their ability to select the fairest models is even weaker, as indicated by lower MRR@1 scores compared to both the baseline and random ranking. Considering that these explanations are often more computationally expensive to generate than evaluating validation set fairness, they are not practically useful as a model fairness indicator. Therefore, we do not recommend explanation-based model selection, especially in decoder-only models. The difference in findings between RQ1 and RQ2 may stem from the fact that debiasing methods can alter model behaviors and thereby affect explanation attributions. As a result, comparing explanations across default and debiased models is less reliable, whereas comparing explanations within the same model remains effective for detecting biased predictions.
@@ 4.4 RQ3: Explanations for Bias Mitigation
While Dimanov et al. (2020) tune hyperparameters based on task accuracy, we search $\alpha\in\{0.01,0.1,1,10,100\}$ using a fairness-balanced metric (the harmonic mean of accuracy and 100–unfairness) on the validation set. Models are selected separately for each fairness criterion. Due to computational cost, we restrict training to single bias types. We exclude DeepLift and KernelSHAP, as they are not easily differentiable and thus cannot be incorporated into model training, and IntGrad, due to its substantial time and memory costs of generating explanations and tracking gradients. Results are averaged over three runs. More implementation details are provided in Appendix A.
@@ Results
In Figure 4, we present race and gender bias mitigation results. For consistency with accuracy, fairness results are reported as $100-\{\text{Disp}_{\text{acc}},\text{Disp}_{\text{fpr}},\text{Disp}_{\text{fnr}},\text{Avg}_{\text{iu}}\}$, so that higher values indicate lower bias. We find that explanation-based bias mitigation effectively improves fairness across multiple metrics. Most notably, it consistently and substantially reduces $\text{Disp}_{\text{fnr}}$ for all bias types. For gender bias, it also yields considerable reductions in $\text{Disp}_{\text{acc}}$, and $\text{Avg}_{\text{iu}}$ is mitigated for race bias. Moreover, as shown in Figure 18, all group fairness disparity metrics decrease for religion bias. The bias mitigation effects are consistent across all models and are also observed on the Jigsaw dataset (see Figures 18, 19, 20, 21 in Appendix E).
At the same time, explanation-based debiasing maintains a good balance between fairness and accuracy. For example, Grad L1 both increases accuracy and reduces $\text{Disp}_{\text{acc}}$, $\text{Disp}_{\text{fnr}}$, and $\text{Avg}_{\text{iu}}$ for gender bias, while most other explanation methods also achieve better $\text{Disp}_{\text{acc}}$ and $\text{Disp}_{\text{fnr}}$ with marginal or no accuracy loss. Our harmonic fairness–accuracy mean results (Figures 22, 23, 24, 25) further confirm this by showing that explanation-based debiasing almost always achieves comparable or higher harmonic means than both default models and traditional debiasing methods.
Among individual explanation methods, attention and attn flow achieve strong debiasing performance on BERT, whereas Occlusion performs best on RoBERTa, though often at the cost of a larger drop in model accuracy. Overall, IxG L2 and attention-based methods provide robust debiasing while maintaining a favorable fairness–accuracy trade-off across bias types, models, and datasets, as reflected in the harmonic mean results. Our findings differ from those of Dimanov et al. (2020), which we attribute to our fairness-based hyperparameter tuning strategy.
@@ 5 Bias Detection in Explanation-Debiased Models
We conduct this analysis using Qwen3 on the same subset for race bias as in the main experiments (see Appendix A for the prompts used).
The fairness correlation differences from default models are shown in Figure 5. We observe that the impact of explanation-based debiasing on fairness correlations depends on both the explanations used for debiasing and those used for bias detection. Some approaches, such as Grad mean / L2, IxG L2, DeepLift mean / L2, Occlusion, and Occlusion abs, are only marginally, or even positively, affected by debiasing. Their fairness correlation scores (see Figure 26 in Appendix F) further indicate that Occlusion- and L2-based methods (except IntGrad L2) remain reliable for revealing bias in explanation-debiased models. In contrast, attention-based explanations experience substantial drops, particularly when the models themselves are debiased using attention-based methods. Similarly, IntGrad-based explanations show a reduced bias detection ability when the debiasing procedure is also gradient-based. Overall, these findings demonstrate that certain input-based explanations remain effective for detecting biased predictions even in explanation-debiased models. Our results are different from those of Dimanov et al. (2020), likely because their analysis focused solely on attribution magnitudes without considering their relationship to fairness metrics.
@@
Existing research suggests that LLMs could identify and correct bias in their own outputs (Bai et al., 2022; Furniturewala et al., 2024). In this section, we compare the bias detection ability of input-based explanations against LLMs’ own judgments under two paradigms: (1) self-reflection, where LLMs are asked to indicate whether their own predictions rely on bias or stereotypes, and (2) self-attribution, where LLMs choose a K-word rationale from the input, which we then examine for the presence of sensitive tokens.
Table 2 shows that self-reflection is highly conservative in flagging bias: only 86 out of 4,000 predictions are labeled as biased, all of which correspond to toxic predictions. The complete absence of non-toxic cases and the extremely low coverage make self-reflection unreliable as a bias detection method due to low recall. Furthermore, predictions marked as biased by the model actually show lower average individual unfairness than those marked non-biased, indicating poor precision as well.
In contrast, LLMs’ self-attributions are less conservative, producing biased / unbiased judgments across both demographic groups and toxicity classes. However, while biased cases identified by rationales show higher average individual unfairness than unbiased ones, they still perform worse than a simple baseline that flags the top 50% of predictions ranked by absolute Grad L2 reliance scores (Grad L2 Binary). Therefore, we conclude that input-based explanations are more reliable than LLM-as-a-Judge for bias detection.
Table 2: Qwen3 results for detecting bias in its own predictions. “Biased / Unbiased” denotes whether an example is judged as biased or unbiased by the LLM through self-reflection or self-attribution. If the judgments are reliable, $\text{Avg}_{\text{iu}}$ should be higher for biased examples than unbiased ones. For self-reflection, fairness correlation cannot be computed because the model labels no non-toxic predictions as biased. Input-based explanations reveal bias more reliably than LLM judgments.
# Biased / Unbiased $\text{Avg}_{\text{iu}}$ (Biased / Unbiased) Fairness Correlation
Self-reflection 86 / 3914 0.065 / 2.59 -
Self-attribution (K=5) 2063 / 1904 3.55 / 1.49 0.104
Self-attribution (K=10) 2176 / 1474 2.93 / 1.56 0.070
@@ 7 Conclusion
In this work, we present the first comprehensive study linking input-based explanations and fairness in hate speech detection. Our experiments show that (1) input-based explanations can effectively identify biased predictions, (2) they are not reliable for selecting fair models, and (3) they can serve as effective supervision signals during training, mitigating bias while preserving a strong balance between fairness and task performance. We further provide practical recommendations on which explanation methods are best suited for bias detection and bias mitigation. Finally, our analyses demonstrate that explanation-based bias detection remains effective in explanation-debiased models, and they outperforms LLM-as-a-Judge in identifying biased predictions.
8 Ethics Statement
@@
This work investigates explainability and fairness in hate speech detection. Despite the diverse experimental setups considered, our findings remain limited by the coverage of tasks, models, datasets, fairness metrics, and identity terms; as such, results may not generalize across groups or domains and could be susceptible to adversarial attacks. We further caution that explanation methods and debiasing techniques cannot fully eliminate residual harms, and that LLM-generated bias judgments are unreliable for bias detection. We hope that our study will contribute to the development of NLP systems that are more transparent, reliable, and fair.
9 Reproducibility Statement
@@
We include full implementation details in the main text and appendix, covering data pre-processing details, model architectures, training procedures, and hyperparameters. We have submitted our code and configuration files as supplementary material to facilitate reproduction during the review process. Upon acceptance, we will open-source our code and scripts for data pre-processing and experiments.
10 Acknowledgements
@@
This work was funded by the DFG project GRK 2853 "Neuroexplicit Models of Language, Vision, and Action" (project number 471607914).
Appendix A Details on Experimental Setups
@@
We use the Captum API [^5] https://captum.ai/api/ to generate explanations, except for Attention, Attn Rollout, and Attn Flow, which we implement ourselves.
@@
Self-Reflection and Self-Attribution
@@
Appendix B Model Performance and Fairness Results

### v2

@@ Results
As a baseline, we report results of using the validation set average individual unfairness as the predictor of test set fairness performance. The results are averaged over six and three random validation set selections for encoder- and decoder-only models, respectively. Results for more models and the Jigsaw dataset are presented in Appendix F.
The results in Figures 3 and 4 highlight the limitations of using explanations for model selection. Although some methods occasionally show high rank correlations (e.g., Grad L2 for race and religion biases in BERT and Occlusion-based methods for gender and religion biases in Qwen3-4B), none of them consistently reach the baseline of using the individual unfairness on the validation set. This limitation is particularly evident in decoder-only models, where the baseline achieves a perfect rank correlation of 1. Similarly, the baseline consistently achieves the highest MRR@1 scores, further showing the limited effectiveness of explanation-based methods in selecting the fairest models. Considering that these explanations are often more computationally expensive to generate than evaluating validation set fairness, they are not practically useful as a model fairness indicator. Therefore, we do not recommend explanation-based model selection, especially in decoder-only models. The difference in findings between RQ1 and RQ2 may stem from the fact that debiasing methods can alter model behaviors and thereby affect explanation attributions. As a result, comparing explanations across default and debiased models is less reliable, whereas comparing explanations within the same model remains effective for detecting biased predictions.
@@ 4.4 RQ3: Explanations for Bias Mitigation
While Dimanov et al. (2020) tune hyperparameters based on task accuracy, we search $\alpha\in\{0.01,0.1,1,10,100\}$ using a fairness-balanced metric (the harmonic mean of accuracy and 100–unfairness) on the validation set [^4] As Occlusion is sensitive to the debiasing strength, we use $\alpha\in\{0.002,0.004,0.006,0.008,0.01\}$.. Models are selected separately for each fairness criterion and results are averaged over three runs. Due to computational cost, we restrict training to single bias types. We exclude DeepLift, DecompX, and KernelSHAP, as they are not easily differentiable and thus cannot be incorporated into model training. Integrated Gradients is substantially more expensive in time and memory for generating explanations and tracking gradients, sp we apply them only to race bias mitigation in BERT and report the results in Table 7 in Appendix G. More implementation details are provided in Appendix B.
@@ Results
In Figure 5, we present race and gender bias mitigation results. For consistency with accuracy, fairness results are reported as $100-\{\text{Disp}_{\text{acc}},\text{Disp}_{\text{fpr}},\text{Disp}_{\text{fnr}},\text{Avg}_{\text{iu}}\}$, so that higher values indicate lower bias. We find that explanation-based bias mitigation effectively improves fairness across multiple metrics. Most notably, it consistently and substantially reduces $\text{Disp}_{\text{fnr}}$ for all bias types. For gender bias, it also yields considerable reductions in $\text{Disp}_{\text{acc}}$, and $\text{Avg}_{\text{iu}}$ is mitigated for race bias. Moreover, as shown in Figure 19, all group fairness disparity metrics decrease for religion bias. The bias mitigation effects are consistent across all models and are also observed on the Jigsaw dataset (see Figures 19, 20, 21, 22 in Appendix G).
At the same time, explanation-based debiasing maintains a good balance between fairness and accuracy. For example, Grad L1 both increases accuracy and reduces $\text{Disp}_{\text{acc}}$, $\text{Disp}_{\text{fnr}}$, and $\text{Avg}_{\text{iu}}$ for gender bias, while most other explanation methods also achieve better $\text{Disp}_{\text{acc}}$ and $\text{Disp}_{\text{fnr}}$ with marginal or no accuracy loss. Our harmonic fairness–accuracy mean results (Figures 23, 24, 25, 26) further confirm this by showing that explanation-based debiasing almost always achieves comparable or higher harmonic means than both default models and traditional debiasing methods.
Among individual explanation methods, attention and attn flow achieve strong debiasing performance on BERT, while IxG L1 and L2 consistently yield a good balance between accuracy and fairness across models. Overall, IxG L2 and attention-based methods provide robust debiasing while maintaining a favorable fairness–accuracy trade-off across bias types, models, and datasets, as reflected in the harmonic mean results. Our findings differ from those of Dimanov et al. (2020), which we attribute to our fairness-based hyperparameter tuning strategy. We further show in Appendix K that combining traditional and explanation-based debiasing methods outperforms either approach in isolation.
@@ 5 Bias Detection in Explanation-Debiased Models
The fairness correlation differences from default models are shown in Figure 6. We observe that the impact of explanation-based debiasing on fairness correlations depends on both the explanations used for debiasing and those used for bias detection. Some approaches, such as Grad mean/L2, IxG L2, DeepLift mean/L2, Occlusion, and Occlusion abs, are only marginally, or even positively, affected by debiasing. Their fairness correlation scores (see Figure 32 in Appendix L) further indicate that Occlusion- and L2-based methods (except IntGrad L2) remain reliable for revealing bias in explanation-debiased models. In contrast, attention-based explanations experience substantial drops, particularly when the models themselves are debiased using attention-based methods. Similarly, IntGrad-based explanations show a reduced bias detection ability when the debiasing procedure is also gradient-based. Overall, these findings demonstrate that certain input-based explanations remain effective for detecting biased predictions even in explanation-debiased models. Our results are different from those of Dimanov et al. (2020), likely because their analysis focused solely on attribution magnitudes without considering their relationship to fairness metrics.
@@
Existing research suggests that LLMs could identify and correct biased model outputs (Bai et al., 2022; Furniturewala et al., 2024). In this section, we compare the bias detection ability of input-based explanations against LLMs’ judgments under two paradigms: (1) LLM decision, where LLMs are asked to indicate whether a model’s prediction rely on bias or stereotypes, and (2) LLM attribution, where LLMs choose a K-word rationale from the input, which we then examine for the presence of sensitive tokens. We conduct this analysis using two LLMs, Qwen3-4B and GPT-OSS-120B, on predictions made by Qwen3-4B on the race subset of Civil Comments (see Appendix B for the prompts used).
Table 2 shows the results of LLM-as-a-judge for bias detection. Under the LLM decision setup, Qwen3-4B is extremely conservative: it flags only 86 out of 4000 predictions as biased, and all of them correspond to toxic predictions. Moreover, the predictions labeled as biased by the model exhibit lower average individual unfairness than those labeled as non-biased, indicating poor precision as well. Under LLM attribution, Qwen3-4B performs slightly better: predictions whose rationales contain sensitive tokens show higher average individual unfairness than those without. However, this still falls short of a simple input-based explanation baseline that flags the top 50% of predictions ranked by absolute Grad L2 reliance scores (Grad L2 Binary). The larger GPT-OSS-120B exhibits improved bias detection ability in the LLM decision setting, but its performance under LLM attribution remains comparable to Qwen3-4B and still substantially worse than input-based explanations. Overall, we conclude that input-based explanations are more reliable than LLM-as-a-judge for bias detection. This finding aligns with the observations of Yang et al. (2025b), who also report that LLM-as-a-judge is unreliable for bias detection.
Table 2: Results of LLM-as-a-judge for bias detection using Qwen3-4B and GPT-OSS-120B. Predictions come from Qwen3-4B on race-related Civil Comments examples. ”Biased/Unbiased” denotes whether an example is judged as biased or unbiased by the LLM through LLM decision or LLM attribution. If the judgments are reliable, $\text{Avg}_{\text{iu}}$ should be higher for biased examples than unbiased ones. For LLM decision with Qwen3-4B, fairness correlation cannot be computed because the model labels no non-toxic predictions as biased. Input-based explanations reveal bias more reliably than LLM-as-a-judge.
LLM Method # Biased/Unbiased $\text{Avg}_{\text{iu}}$ (Biased/Unbiased) Fairness Correlation
Qwen3-4B LLM decision 86/3914 0.065/2.59 -
LLM attribution (K=5) 2063/1904 3.55/1.49 0.104
LLM attribution (K=10) 2176/1474 2.93/1.56 0.070
GPT-OSS-120B LLM decision 399/3601 4.42/2.35 0.051
LLM attribution (K=5) 2153/1843 3.33/1.65 0.092
LLM attribution (K=10) 2729/1238 2.88/1.74 0.063
@@ 7 Conclusion
In this work, we present the first comprehensive study linking input-based explanations and fairness in hate speech detection. Our experiments show that (1) input-based explanations can effectively identify biased predictions, (2) they are not reliable for selecting fair models, and (3) they can serve as effective supervision signals during training, mitigating bias while preserving a strong balance between fairness and task performance. We further provide practical recommendations on which explanation methods are best suited for bias detection and bias mitigation. Finally, our analyses demonstrate that explanation-based bias detection remains effective in explanation-debiased models, and they outperforms LLM-as-a-judge in identifying biased predictions [^5] Limitations and future directions are discussed in Appendix M. .
Ethics Statement
@@
This work investigates explainability and fairness in hate speech detection. Despite the diverse experimental setups explored and the additional generalization tests in Appendix H, the findings are still constrained by the specific configurations considered here. As such, the results may not fully generalize across demographic groups, domains, or tasks, and they may remain vulnerable to adversarial manipulation. We further caution that explanation methods and debiasing techniques cannot fully eliminate residual harms, and that LLM-generated bias judgments are unreliable for bias detection. We hope that our study will contribute to the development of NLP systems that are more transparent, reliable, and fair.
Reproducibility Statement
@@
We include full implementation details in the main text and appendix, covering data pre-processing details, model architectures, training procedures, and hyperparameters. Our code and scripts for data pre-processing and experiments are available at https://github.com/Ewanwong/fairness_x_explainability.
Acknowledgements
@@
This work was funded by the DFG project GRK 2853 ”Neuroexplicit Models of Language, Vision, and Action” (project number 471607914).
We gratefully acknowledge the computing time made available to them on the high-performance computer at the NHR Center of TU Dresden. This center is jointly supported by the Federal Ministry of Research, Technology and Space of Germany and the state governments participating in the NHR (www.nhr-verein.de/unsere-partner).
Mayank Jobanputra is supported by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project ID 389792660-TRR 248 (Foundations of Perspicuous Software Systems).
We also sincerely thank the reviewers and area chairs for their exceptionally detailed and constructive feedback.
Appendix A Fairness and Explainability in Hate Speech Detection
To better motivate our focus on fairness and explainability in the hate speech detection task, we provide additional background in this section. We begin by clarifying our definitions of hate speech and social bias, then review relevant work on fairness and explainability in hate speech detection. Finally, we explain why our study specifically focuses on input-based explanations.
Hate Speech
We follow (Fortuna and Nunes, 2018) in defining hate speech as a specific form of abusive or toxic language that targets and attacks protected or identifiable social groups. Under this view, hate speech is a subset of abusive language. This definition is consistent with widely adopted formulations in prior work on hate speech detection (e.g., Nobata et al., 2016; Davidson et al., 2017). Because our study focuses on fairness and, in particular, analyzes model behavior on examples involving specific social groups (race, gender, religion), this standard definition of hate speech aligns well with the scope and goals of our work. We still use the toxic vs. non-toxic labels following the terminology used in the Civil Comments (Borkan et al., 2019) and Jigsaw (cjadams et al., 2019) datasets. Although these datasets include multiple subtypes of abusive content, they group them under the broader notion of toxicity.
Social Bias
Following the conceptualization of (Blodgett et al., 2020), we define social bias as the presence of stereotypical associations between social groups and certain attributes, as well as disparities in how these groups are treated as a result. Such biases can lead to both representational harms (e.g., demeaning or misrepresenting targeted groups) and allocational harms (e.g., unfair distribution of opportunities or resources). Given the potential for NLP systems to reproduce or amplify these harms, and their growing influence in everyday life, it is essential to detect and mitigate social bias in these models.
Social Bias in Hate Speech Detection
Social bias has been widely documented in hate speech detection systems. Dixon et al. (2018) showed that training data often contain uneven distributions of identity terms and stereotypical associations, which in turn propagates bias into downstream models. Subsequent studies revealed multiple dimensions of such disparities: Sap et al. (2019) demonstrated systematic dialectal prejudice against African-American English (AAE), while Park et al. (2018) reported significant performance gaps across gendered identities. Garg et al. (2019) further found that models frequently assign different toxicity labels to otherwise identical content when only the referenced social group is varied. Sahoo et al. (2022) expanded the scope of this line of study by curating the ToxicBias dataset and examining bias across a broader set of social categories. More recently, Roy et al. (2023) found that LLMs exhibit similar bias in hate speech detection. Together, these studies underscore the persistence and multifaceted nature of social bias in hate speech detection.
To address such biases, a rich line of work has proposed mitigation techniques at different stages of the modeling pipeline. Pre-processing methods include debiasing word embeddings to reduce spurious associations between identity terms and toxicity (Park et al., 2018), re-sampling or re-weighting examples to obtain more balanced label distributions across identity groups (Dixon et al., 2018), and counterfactual data augmentation (Park et al., 2018; Garg et al., 2019). In-processing approaches mostly modify the training objective, for instance by adding fairness-aware regularizers that penalize correlations between identity terms and toxic predictions (Garg et al., 2019; Davani et al., 2020; Attanasio et al., 2022; Schäfer et al., 2024). Post-processing methods adjust model outputs without retraining (Wang and Demberg, 2024b): threshold adjustment per group has been used to trade off subgroup false positive and false negative rates and reduce disparities (Dixon et al., 2018), while Mamta et al. (2024) identify neurons associated with biased behavior and prune or edit them to improve fairness.
Despite substantial progress on identifying and mitigating social bias in hate speech detection, relatively little work has systematically explored whether model explanations can be leveraged to detect or reduce such biases.
Explainability in Hate Speech Detection
In parallel, there is a growing line of work on input-based explanations for hate speech detection. HateXplain (Mathew et al., 2021) introduces a benchmark with human-annotated rationales and shows that models trained with rationale supervision improve both interpretability and reduce unintended bias towards target communities. Building on this, Kim et al. (2022) and Saha et al. (2023) train models to jointly predict human rationales and toxicity labels, leading to more robust and explainable hate speech detection systems. More recent work further leverages LLM-generated rationales to supervise hate speech classifiers, achieving improved performance and interpretability (Nirmal et al., 2024).
However, while existing efforts focus primarily on improving hate speech detection performance, relatively little work examines whether and how input-based explanations can be systematically leveraged to improve fairness in hate speech detection models. Since fairness and explainability have both been extensively studied in this task, hate speech detection serves as an ideal setting for a thorough empirical examination of how these two dimensions interact in NLP models.
Input-Based Explanations
We focus on input-based explanations because they offer the most direct view into which parts of the input influence a model’s prediction (Wang and Yin, 2021), and they have long been regarded as central tools for fairness auditing in ML (Balkir et al., 2022; Deck et al., 2024). Their methodological diversity also makes them an ideal testbed for our study, enabling a comprehensive examination of whether and how explanations can improve fairness (Lyu et al., 2024). In addition, both automated and human-centric metrics for evaluating explanation properties (e.g., faithfulness, interpretability) are well established (DeYoung et al., 2020; Jacovi and Goldberg, 2020; Lage et al., 2019). This allows us to analyze how these properties relate to an explanation method’s (in)effectiveness in fairness-related tasks. Finally, input-based explanations are often mandated by laws, such as the EU Artificial Intelligence Act, making it practically important to understand how their use interacts with fairness considerations.
Appendix B Details on Experimental Setups
@@
(Now, please analyze the following sentence and provide your
answer. Answer only Yes or No.)
@@
We implement Attention, Attention Rollout, Attention Flow, and Progressive Inference ourselves, and use the official codebase from the DecompX paper for its implementation. For other explanation methods, we generate them using the Captum API [^8] https://captum.ai/api/ Single-pass Progressive Inference is used for simplicity.
@@
LLM decision and LLM attribution
@@
Appendix C Explanation Efficiency
Table 4 reports the time and GPU memory costs for each explanation method. Most post-hoc explanation methods are lightweight when applied to BERT, whereas IntGrad, Occlusion, and KernelSHAP require substantially more time and computational resources when generating explanations for LLMs.
Table 4: Computational costs per example for generating explanations across 200 instances on BERT and Qwen3-4B. Results are computed on the race subset of the Civil Comments dataset using a batch size of 1 and are averaged over three runs. All methods are run on a single 80-GB H100 GPU, except Integrated Gradients, which uses two H100 GPUs with gradient checkpointing to reduce memory usage. Because explanation methods within the same family incur similar computational costs, we report each family only once.
BERT Qwen3-4B
Method Time (s/example) Memory (GB) Method Time (s/example) Memory (GB)
Attention 0.027 0.529 Attention 0.112 16.598
Grad 0.026 0.603 Grad 0.237 19.631
IxG 0.025 0.603 IxG 0.236 19.631
IntGrad 0.064 7.010 IntGrad 1.784 101.694
DeepLift 0.027 0.748 DeepLift 0.323 23.530
Occlusion 0.330 0.508 Occlusion 4.204 15.639
KernelSHAP 0.138 0.508 KernelSHAP 1.374 20.013
DecompX 0.074 7.222 ProgInfer 0.068 15.810
Appendix D Model Performance and Fairness Results